Session: May 1997 (session 2)
Frequency: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Brand at Thu Jun 5 18:45:21 METDST 1997) Success
During much of the experiment it was raining About 2 hrs into the experiment there is a note in the log about strong interference near the IF bandwidth.

Onsala (A. Polatidis at Thu Jun 5 20:51:58 METDST 1997) Success
No known problems

Effelsberg (dgraham at Fri Jun 6 11:27:05 METDST 1997) Success

Jodrell (Alastair Gunn at Mon Jun 9 10:18:52 METDST 1997) Minor failures
Engineers installed a secondary focus receiver at primary focus, hence polarisations are swapped. This has been corrected for subsequent experiments. No other known problems.

Westerbork (tony foley at Mon Jun 9 14:35:55 METDST 1997) Minor failures
missed passes 1 to 3 on the second tape (~1200-1311UT)

Noto (C.Trigilio at Mon Jun 9 16:07:18 METDST 1997) Minor failures
Right circular polarisation not observed because of a problem in the front end. No other problem.

Torun (K.Borkowski at Tue Jun 10 11:31:40 METDST 1997) Minor failures
Until 13:20 signals recorded with reversed polarisation. 12:40 to 12:48 break in observations.

Urumqi (Li Suqin at Thu Jun 12 07:22:58 METDST 1997) Minor failures
It is claudy , the colck offset =5.7+-0.5us. From 06:35:00 to 07:41:32 did not take the test. schedule 08:10:00 was late start at 08:10:50.

Simeiz (A.Volvatch/LIG at Wed Jun 18 08:53:12 METDST 1997) Success
No problem reported, time offset 04 usec. Tapes shipped on 17.06.97.

Medicina (Fredrik T Rantakyro at Mon Jul 21 11:19:08 METDST 1997) Minor failures
The ANTAB file for the experiment can be found on VLBEER in the MAY97 directory with the name C97C2A.ANTAB. The file is ready to be used with the AIPS tasks ANTAB/APCAL. Read the ANTAB file for any specific comments regarding this experiment.

MPIfR (Scott Aaron at Fri Feb 6 11:48:08 MET 1998) Minor failures
Correlation completed. Only upper sidebands were correlated. Jodrell has no data on the reverse passes.